January 7th 2022

From Hall A Wiki
Jump to: navigation, search
  • Updates
    • HCal
      • LED timing shift
    • Shower
      • will try to take cosmics this evening
      • HV for running ?
      • cosmics trigger
    • will restore the BB low trigger
    • script : jmirror being set up by Bob, checking with Dave Lawrence in Hall D, and Sergei
    • helicity : David will try the helicity
    • CDet
      • DAQ restarted after outage
      • will swap detector around next week
    • GEM
      • working on shielding 4 th chamber UV, chamber working well and being installed
      • GRINCH
    • Timing shift :
      • Bryan did not see any issues with TI / TS setting
      • Bryan want to add informations from TI and TS in
      • Bryan working on automatically disabling bad APVs so run continues without hanging DAQ